Oscar's dog house is shaped like a triangle tent. The slanted sides are both 5 feet long and the bottom of the house is 6 feet across. What is the height of his dog house, in feet, at its tallest point? (Draw and think)
Answer by richwmiller(17219)

isosceles triangle 5,5,6
drop a line from vertex to midpoint of base 6 for the height
now we have triangle base 3 hypotenuse 5
3^2+h^2=5^2
9+h^2=25
h^2=16
h=4 ft
3,4,5 right triangle
The dog house consists of two 3,4,5 right triangle